# YOLO11 and Vision Transformers based 3D Pose Estimation of Immature Green Fruits in Commercial Apple Orchards for Robotic Thinning

### v1

@@ 3 Methods
The overall methodology of this study is summarized in Figure 2. In this study, RGB images were collected using a robotic system, as depicted in Figure 2. This figure comprehensively illustrates the workflow, encompassing data collection, preprocessing, and the deployment of deep learning models for pose estimation. The images obtained were manually annotated through extensive human labor, which involved the precise delineation of bounding boxes and pose lines on immature green fruits, crucial for the subsequent model training phase. These annotated datasets were employed to train various configurations of the YOLO11 and YOLOv8 models, with the comparative performance analysis of each configuration presented in the middle part of Figure 2. The aim was to ascertain the most efficacious model configuration based on precision and inference speed for applications in robotic agricultural settings. The optimal model configuration was subsequently utilized for the transformation of RGB images into RGB-D point clouds. This transformation was facilitated by two advanced vision transformers: the Dense Prediction Transformer (DPT) and Depth Anything V2. The conversion from RGB to RGB-D data was critical for the validation process, detailed on the right side of Figure 2. The estimated poses derived from the YOLO11 model were validated against actual field measurements obtained using a digital caliper. This validation process involved measuring the major axis length of each fruitlet in the RGB-D point clouds and comparing these measurements with the corresponding field data to ensure the accuracy and reliability of the pose estimations.
The methodologies employed in this study are detailed in the following five sections. Initially, the Data Acquisition section discusses how images were collected. This is followed by an overview and training of the YOLO11 models, as described in the YOLO11 Overview and Model Training section. The performance of these models is then evaluated in the Detection and Pose Estimation Performance Evaluation section. Subsequently, the Vision Transformers for RGB to RGB-D Mapping section is divided into two subsections: Depth Prediction Transformer (DPT) and Depth Anything V2. Finally, the Pose Validation in a 3D Environment section details how the accuracy of our pose estimates was assessed against actual measurements.
Figure 2: Outlining the methodology from data collection via a robotic system, through deep learning model training and validation, to the precise pose estimation of immature green fruitlets in commercial orchard environment.
@@ 3.1 Data Acquisition and Preparation
Data acquisition was carried out using a robotic platform that was equipped with an Intel RealSense D435i camera, mounted on a UR5e robotic arm as depicted in Figure 2. This setup facilitated the capture of RGB images of immature fruitlets in a commercial "Scifresh" apple orchard located in Prosser, Washington State, USA. The images were collected prior to thinning in May 2024, with the appropriate timing for fruitlet thinning determined through regular monitoring of the commercial apple orchard and ongoing communication with growers and laborers. The orchard was planted with tree row spacing of 10 feet apart and a spacing of 3 feet between trees.
@@
A total of 1147 images were collected using Intel RealSense D435i camera as shown in Figure 2. This camera features a depth-sensing system utilizing active infrared (IR) stereo vision and an inertial measurement unit (IMU). The camera’s depth sensor operates using structured light technology, which employs a pattern projector to create disparities between stereo images captured by two IR cameras. With a resolution of 1280 × 720 pixels, the camera can capture depth information up to a range of 10 meters. It supports a frame rate up to 90 frames per second (fps), offering a 69.4° horizontal field of view (HFOV) and 42.5° vertical field of view (VFOV). Additionally, the 6-axis IMU provides precise orientation data, enhancing the alignment of depth data and scene understanding.
@@
After the image collection phase, the images were then preprocessed to enhance their quality. The preprocessing was followed by a manual annotation process conducted by experts. Each image was annotated to identify crucial keypoints on the immature fruits, specifically the calyx and peduncle, as illustrated in Figures 2 and 3. A bounding box and a line were used to denote the object and its pose, respectively. This extensive manual annotation process required continuous effort over several days by multiple individuals. It was vital for generating accurate data essential for training deep learning models. After this labor-intensive phase, the dataset was carefully divided into 918 images for training, 114 for validation, and 115 for testing, adhering to an 8:1:1 train:validate ratio for data preparation. This organized dataset was then exported using Roboflow, ensuring structured access for model development.
@@
Figure 3: YOLO11 Architecture diagram for immature green fruit detection and pose estimation
3.2 YOLO11 Overview and Model Training
@@
YOLO11 (Glenn Jocher, Ultralytics, 2024, Link: https://docs.ultralytics.com/models/yolo11/ ) (Figure 3) represents the latest advancement in the YOLO family, building on the strengths of its predecessors while introducing innovative features and optimizations aimed at enhancing performance across various computer vision tasks. Specifically, YOLO11 architecture optimized feature extraction capabilities as depicted in Figure 3, enabling it to capture intricate details in images. This model supports a range of applications, including real time object detection, instance segmentation, pose estimation, and which allows it to accurately identify objects regardless of their orientation, scale, or size, making it versatile for industries like agriculture and surveillance. The YOLO11 model utilizes enhanced training techniques that have led to improved results on benchmark datasets. Notably, YOLO11m achieved a mean Average Precision (mAP) score of 95.0% on the COCO dataset while utilizing 22% fewer parameters compared to YOLOv8m, demonstrating greater efficiency without compromising accuracy. With an average inference speed that is 2% faster than YOLOv10, YOLO11 is optimized for real-time applications, ensuring quick processing even in demanding environments. These specifications position YOLO11 as a powerful tool for advancing AI applications, particularly in sectors where rapid and accurate image analysis is crucial.
@@
The training process for the YOLO11 models was structured to optimize performance across various computer vision tasks. The model was trained with a batch size of 8 and a uniform image resolution of 640x640 pixels. The training utilized an automatic optimizer and implemented a total of 100 epochs, with a patience setting of 100 to avoid overfitting. Key hyperparameters included an initial learning rate of 0.01, momentum at 0.937, and weight decay set at 0.0005. A warmup phase of 3 epochs was integrated to stabilize training parameters, alongside specific loss settings including box loss of 7.5 and class loss of 0.5. Image augmentation techniques used were flipping, translation, and a unique mosaic approach with a probability of 1.0. Moreover, the model was designed to operate with advanced features like overlap masks and a dynamic workspace of 4. In this work, pre-initialized weights were not used, ensuring that the models were trained to adapt solely based on the synthesized dataset, thus providing a fresh foundation to assess the effectiveness of the training and data augmentation strategies employed. The training process was conducted on a high-performance computing workstation, which was crucial for managing the substantial computational demands of the study.
@@
The workstation was equipped with an Intel Xeon(R) W-2155 CPU, which features a base clock speed of 3.30 GHz and 20 cores, providing the necessary processing power for intensive data processing tasks. For graphics processing and machine learning computations, the system utilized NVIDIA Corporation GP102 [TITAN Xp] graphics cards, which are well-suited for complex image processing tasks. Additionally, the workstation boasted a large storage capacity of 7.0 TB, allowing for extensive data management and analysis. It operated under Ubuntu 20.04.6 LTS, a robust and stable 64-bit operating system. The system’s graphical interface was managed by GNOME version 3.36.8, with X11 serving as its windowing system, ensuring a stable and efficient environment for conducting the computational analyses required for training the YOLO11 configurations.
@@
After the YOLO11 experiments, its predecessor, YOLOv8, along with its five configurations (developed by the same company, Ultralytics), were thoroughly evaluated. Each configuration of YOLOv8 was tested under identical hyperparameter settings and using the same workstation to ensure a controlled and fair comparison.
@@ 3.3 Performance Evaluation for Detection and Pose Estimation
The detection and pose estimation performance of the YOLO11 algorithm for immature green apples was evaluated using key metrics: precision, recall, and mean Average Precision at 50% Intersection over Union (mAP@50). These metrics are calculated based on the following terms:
@@
These metrics, delineated through Equations 1, 2, and 3, provide a comprehensive view of the model’s performance in detecting and estimating the pose of immature green fruits, highlighting the effectiveness and accuracy of the detection algorithms.
Additionally, the evaluation of image processing speeds for green fruit pose detection was systematically segmented into three phases: pre-processing, inference, and post-processing. These phases encompass the complete cycle from initial image manipulation before detection, through the detection process itself, to the final steps following the detection output. Each phase is critical in understanding the overall efficiency and responsiveness of the pose detection system.
@@
  • Inference: At this core phase, the model processes the image to detect and estimate the pose of the fruits. The speed of this stage is crucial for real-time applications and is quantified by the time taken to process a single image, as shown in Equation 4:
@@ 3.4 Vision Transformers for RGB to RGB-D Mapping
The two Vision Transformers, DPT and Depth Anything V2, are explored in this section. Detailed workflows for each model are illustrated in Figure 4, and the specific processes for DPT and Depth Anything are further elaborated in Figures 6 and 5.
DPT, designed primarily for depth estimation, utilizes a transformer architecture that excels in processing high-resolution RGB images to generate depth-enhanced RGB-D data efficiently. This model leverages pre-trained weights and a sophisticated feature extraction process to offer precise depth information, crucial for accurate 3D mapping of objects. In contrast, Depth Anything V2 extends the capabilities of depth prediction to various scenes by adapting to different datasets without extensive retraining. It uses a lightweight transformer model that supports faster computation while maintaining high accuracy. This model is particularly effective in generating detailed depth maps from RGB images, which are instrumental in calculating the exact dimensions and orientations needed for robotic thinning operations in orchards.
Figure 4: Illustrating the methodology for depth mapping and validation using Vision Transformers in the study. It details the process from RGB image input through two transformer models, DPT and Depth Anything V2, for generating RGB-D data, resulting in high-quality 3D point cloud creation and validation against ground truth measurements.
Figure 5: Workflow of the Dense Prediction Transformer (DPT) for converting RGB images to RGB-D data. This diagram shows the detailed process involved in transforming standard RGB images into depth-inclusive RGB-D formats using the DPT model.
Figure 6: Workflow of Depth Anything V2, detailing each step from image collection to depth map generation. This diagram illustrates how Depth Anything V2 processes input images to produce depth maps, contributing to 3D model reconstruction.
@@ 3.4.1 Depth Prediction Transformer (DPT)
The DPT model utilizes a ViT as its backbone, which processes images in a sequence of patches. Unlike standard transformers, DPT assembles tokens from different stages of ViT into multi-resolution feature maps, which are then progressively fused using a convolutional decoder to produce full-resolution predictions. This method ensures that the model captures both fine details and global context effectively. The model operates on a high constant resolution throughout the process, enhancing its ability to generate detailed and coherent predictions across the entire image.
In this study, the DPT was directly applied to high-resolution RGB images collected from commercial apple orchards to generate detailed RGB-D (depth) data. The depth maps produced by DPT provided the 3D spatial context necessary for validating the pose estimation performed by the YOLO11 algorithm. Specifically, YOLO11’s pose estimations of the major axes of immature green fruits were cross-verified against measurements obtained from 3D point clouds generated by DPT. The actual diameters of the fruitlets were measured in the field using digital calipers to provide ground truth validation. This integration of DPT not only enhanced the accuracy of pose estimation but also demonstrated the potential of transformers in automating and refining agricultural operations through precise 3D mapping and measurement capabilities.
The entire workflow, as depicted in the accompanying diagram (refer to Figures 6 and 5) , was meticulously designed to ensure a streamlined process. Initially, high-resolution RGB images were captured using specialized cameras mounted on a robotic platform. These images were then preprocessed to enhance quality before being input into the DPT model. The model processed these images through several stages, including patch extraction, transformer encoding, and feature map creation, culminating in the generation of detailed RGB-D outputs. The subsequent steps involved using these outputs to validate the pose estimations made by the YOLO11 algorithm, focusing on the accuracy of keypoint detection within the regions of interest. Each step of the process was critical in achieving a high degree of precision in the 3D spatial analysis, ultimately facilitating a robust assessment of the pose estimation techniques used. The methodology outlined ensured that the study leveraged advanced imaging and deep learning techniques to address the challenges posed by manual labor in fruitlet thinning operations.
@@ 3.4.2 Depth Anything V2
Depth Anything V2 is an advanced monocular depth estimation model designed to convert standard RGB images into depth-mapped (RGB-D) images. This model employs a cutting-edge neural network architecture that integrates depth estimation directly into the image processing pipeline, enabling it to predict the depth information from a single image with high precision. The model utilizes a fully convolutional network, leveraging the power of deep learning to interpret and transform visual data into spatial mappings. The Depth Anything V2 architecture is characterized by its depth prediction capabilities, which are crucial for applications requiring detailed three-dimensional understanding from two-dimensional data. The model processes images through several layers, where each layer extracts and refines features at different scales. This method allows the model to capture both high-level details and fine-grained elements, making it exceptionally good at handling the complexities of natural scenes. It uses a series of encoder-decoder networks, where the encoder compresses the input image into a feature-rich representation, and the decoder expands these features back into an image format, assigning depth values to each pixel. This is enhanced by skip connections that help preserve spatial details that are often lost during down-sampling. The output is a dense depth map that closely aligns with the real-world geometry of the photographed scene.
In this study, the Depth Anything V2 model was utilized to significantly enhance the pose estimation of fruitlets, which is crucial before the thinning process in agricultural operations. This methodology was systematically executed in a sequence of defined steps, as depicted in the associated diagram (refer to Figures 6 and 5). Initially, high-resolution RGB images of immature green fruitlets were captured using a specialized camera setup in commercial orchards, as outlined in the "Fruitlet Image Collection" step. These images were then subjected to preprocessing to optimize quality for further analysis, a process detailed in the "Image Preprocessing" step. Following this, the images were inputted into the Depth Anything V2 model, starting the depth estimation process, as indicated in the "Depth Estimation Initialization" step. Subsequent steps involved extracting features from the images ("Feature Extraction"), compressing these features via an encoder to generate a dense representation ("Encoder Processing"), and then expanding these representations through a decoder to form detailed depth maps ("Decoder Processing"). Integration of skip connections was employed to preserve crucial spatial details, a key feature described in the "Skip Connection Integration" step. The depth maps generated were used to synthesize 3D spatial data about the fruitlets ("3D Information Synthesis"), which enhanced the YOLO11 algorithm’s ability to accurately localize and measure each fruit in three dimensions. This critical phase is captured under the "Pose Estimation Integration" step. The synthesized depth data provided a detailed understanding of the spatial arrangement of fruitlets, crucial for identifying precise engagement points for robotic arms during thinning.

### v2

@@ 3 Methods
The overall methodology of this study is summarized in Figure 2. As show in the figure, the workflow consists of data collection, pre-processing, and the deployment of deep learning models for pose estimation. A dataset consisting of RGB and Depth images was collected from a commercial apple orchard using a robotic platform. The images obtained were manually annotated, which involved precise delineation of bounding boxes and estimation of pose lines on immature green fruits, crucial for the subsequent model training phase. These annotated datasets were employed to train various configurations of the YOLO11 and YOLOv8 models for comparative assessment of their performances. The goal of this step was to identify the most effective model configuration for fruitlet pose estimation based on the precision and inference speed achieved. The optimal model identified from this comparative assessment was then used to transform RGB images into RGB-D point clouds. This process to generate 3D point clouds utilized monocular images and two state-of-the-art Vision Transformers: Depth Anything and Dense Prediction Transformer. These tools mapped the RGB properties of the images to simulate depth, creating depth-enhanced RGB-D data necessary for accurate 3D representation and enhanced pose estimation. The conversion from RGB to RGB-D data was also critical for the validation process as depicted in the right side of Figure 2. The poses estimated by the YOLO11 model were validated against actual field measurements obtained using a digital caliper. This validation process involved estimating the major axis length of each fruitlet in the RGB-D point clouds and comparing these measurements with the corresponding field measurements to ensure the accuracy and reliability of the pose estimation models. In the following several sub-sections, these steps of the research methodology, including ”Data Acquisition”, ”Model Training” and ”Field Validation” will be described in more details.
Figure 2: Outlining of the research the methodology that includes various steps including data collection (using a robotic platform), deep learning model training and validation, and precise pose estimation of green fruitlets.
@@ 3.1 Data Acquisition and Preparation
Data acquisition was carried out using a robotic platform that was equipped with an Intel RGB-D camera, mounted on a UR5e robotic arm as depicted in Figure 2. This setup facilitated the capture of RGB images of immature fruitlets in a commercial ”Scifresh” apple orchard located in Prosser, Washington State, USA. The images were collected prior to fruitlet thinning in May 2024, with the appropriate timing for fruitlet thinning determined through regular monitoring of the commercial apple orchard and ongoing communication with growers and farm workers. The orchard used was planted in 2008 with tree row spacing of 3 meter and an intra-row tree spacing of 1 meter.
@@
In this study, a total of 1,147 images were collected using Intel RealSense D435i camera. This camera features a depth-sensing system utilizing active infrared (IR) stereo vision and an inertial measurement unit (IMU). The camera’s depth sensor operates using structured light technology, which employs a pattern projector to create disparities between stereo images captured by two IR cameras. The 3D sensor of this camera has a resolution of 1280 × 720 pixels and can capture depth information up to a range of 10 m. The camera supports a frame rate up to 90 frames per second (fps), and has a horizontal field of view (HFOV) of 69.4° and vertical field of view (VFOV) of 42.5°. Additionally, the 6-axis IMU provides precise orientation data, enhancing the alignment of depth data and scene understanding.
@@
The pre-processed images manually annotated in roboflow (Roboflow, Des Moines, IOWA) manually to create the dataset for model training and testing. The annotation process included identification of y crucial keypoints on immature fruitlets, specifically the calyx and peduncle, as illustrated in Figures 2 and 3. A bounding box and a line were then used to denote the object and its pose, respectively. This was a labor-intensive annotation process, which was vital for generating accurate object labels essential for training deep learning models. Annotated dataset was then divided into 80% (918 images) for training, 10% (114 images) for validation, and 10% (115) for testing. This organized dataset was then exported using Roboflow, ensuring structured access for model development and training.
@@
Figure 3: YOLO11 Architecture used for immature green fruit detection
3.2 YOLO11 Model Overview and Training for Pose Estimation
@@
In this study, we employed the YOLO11 (Glenn Jocher 2024 Jocher and Qiu (2024)) model for its advanced capabilities in detecting and estimating the pose of immature green fruits in orchard environments. YOLO11 (Figure 3) represents the latest advancement in the YOLO series, building on the strengths of its predecessors with innovative features and optimizations that enhance accuracy and processing speed, making it ideal for real-time applications in agricultural environments where rapid and precise object detection is crucial. Specifically, the YOLO11 architecture has optimized feature extraction capabilities as depicted in Figure 3, enabling it to capture fine details in images. This model supports a wide range of applications, including real-time object detection, instance segmentation, and pose estimation. It is capable of accurately identifying objects irrespective of their orientation, scale, or size, making it highly versatile for use in industries such as agriculture and surveillance. YOLO11 incorporates advanced training techniques that have led to significant improvements on benchmark datasets. For example, the YOLO11m variant achieved a higher mean Average Precision (mAP) on the COCO dataset while utilizing 22% fewer parameters than YOLOv8m, showcasing enhanced efficiency without sacrificing accuracy. Designed with an inference speed that surpasses earlier models, YOLO11 is optimized for real-time applications, ensuring fast processing even in demanding environments.
@@
The training process for the YOLO11 models was structured to optimize performance across various computer vision tasks. The model was trained with a batch size of 8 and a uniform image resolution of 640x640 pixels. The training process utilized an automatic optimizer and implemented a total of 100 epochs, with a patience setting of 100 to avoid over-fitting. Key hyperparameters optimzied included an initial learning rate of 0.01, momentum at 0.937, and weight decay set at 0.0005. A warmup phase of 3 epochs was integrated to stabilize training parameters, alongside specific loss settings including box loss of 7.5 and class loss of 0.5. The default model image augmentation techniques used were flipping, translation, and a unique mosaic approach with a probability of 1.0. Moreover, the model was designed to operate with advanced features like overlap masks and a dynamic workspace of 4. In this work, pre-initialized weights were not used, ensuring that the models were trained to adapt solely based on the dataset collected in this study, thus providing a fresh foundation to assess the effectiveness of the training and data augmentation strategies employed.
@@
To meet the significant computational needs of this study, the training was conducted on a high-performance computing system. This system featured an Intel Xeon(R) W-2155 CPU with a base clock speed of 3.30 GHz and 20 cores. For graphics processing and machine learning tasks, the system was equipped with NVIDIA Corporation GP102 [TITAN Xp] graphics cards, adept at handling complex image processing tasks. The workstation also boasted a large storage capacity of 7.0 TB, facilitating extensive data management and analysis. It ran on Ubuntu 20.04.6 LTS, a robust and stable 64-bit operating system. The graphical interface of the system was managed by GNOME version 3.36.8, and X11 was used as its windowing system, creating a stable and efficient environment for the computational analyses required to train the YOLO11 configurations.
@@
After the YOLO11 training was completed, its predecessor, YOLOv8, along with its five configurations (developed by the same company, Ultralytics), were thoroughly evaluated. Each configuration of YOLOv8 was tested under identical hyperparameter settings and using the same workstation to ensure a controlled and fair comparison.
@@ 3.3 Performance Evaluation for Detection and Pose Estimation
The performance of YOLO11 model in detecting immature green apples and estimating their pose was evaluated using precision, recall, and mean Average Precision at 50% Intersection over Union (mAP@50). These metrics are further defined as follows:
  • 1.
@@
These metrics (1, 2, and 3) provide a comprehensive view of the model’s performance in detecting and estimating the pose of immature green fruits, highlighting the effectiveness and accuracy of the models used.
Additionally, the evaluation of image processing speed for green fruit pose detection was systematically divided into three phases: pre-processing, inference, and post-processing. These phases encompass the complete cycle from initial image manipulation before detection, through the detection process itself, to the final steps following the detection. Each phase is critical in understanding the overall efficiency and responsiveness of the pose detection system.
  • 1.
@@
  • 2.
    Inference: At this core phase, the model processes image to detect fruit and estimate their pose. The speed of this stage is crucial for real-time applications and is quantified by the time taken to process a single image, which is given by 4:
@@
Based on these measures, the best-performing YOLO11 configuration was selected from all tested variants of YOLO11 and YOLOv8 pose detection models. The 2D Pose estimation of immature green apples from the selected model configuration were used to implement two Vision Transformer models; DPT and Depth Anything V2. These models transformed standard RGB images into enriched RGB-D point clouds, essential for advancing detailed 3D mapping capabilities. This step was pivotal in integrating precise depth information with RGB image, which is expected to enhance the robotic technologies for effective thinning operations in commercial apple orchards.
@@ 3.4 Vision Transformers for RGB to RGB-D Mapping
The two Vision Transformers, DPT and Depth Anything V2, are explored in this section for point cloud generation based on RGB image. Overall block-diagram for this step is presented in Figure 5a whereas detailed workflow for each model is illustrated respectively in Figures 5a, and 5b.
DPT, designed primarily for depth estimation, utilizes a transformer architecture that is effective in processing high-resolution RGB images to generate depth-enhanced RGB-D data. This model leverages pre-trained weights and a sophisticated feature extraction process to estimate precise depth information, which is essential for accurate 3D mapping of objects. This model uses a lightweight transformer model that supports faster computation while maintaining high accuracy. The model is particularly effective in generating detailed depth maps from RGB images, which are essential in calculating the dimensions and orientations of objects (e.g., size, volume, orientation , distance, texture, segmentation, and camera perspectiveetc) needed for robotic thinning operations in orchards.
Figure 4: Flowchart of the methodology used for depth mapping and validation using two Vision Transformer models. RGB images of apple trees with immature green apples (fruitlets), initially processed using YOLO11n for pose estimation, were transformed into RGB-D data by Dense Prediction Transformer (DPT) and Depth Anything V2. The depth maps generated were then used to create precise 3D point clouds, which were validated against ground truth measurements.
Figure 5: (a) Workflow of the Dense Prediction Transformer (DPT) for generating 3D point clouds. DPT processes RGB images, originally captured using a consumer-grade camera, to produce RGB-D data, facilitating the creation of detailed 3D point clouds; (b) Workflow of Depth Anything V2, detailing each step from RGB image processing to depth map generation and subsequent 3D point cloud generation
@@ 3.4.1 Depth Prediction Transformer (DPT)
The DPT model utilizes a Vision Transformer (ViT) as its backbone, which processes images in a sequence of patches. Unlike standard transformers, DPT assembles tokens from different stages of ViT into multi-resolution feature maps, which are then progressively fused using a convolutional decoder to produce full-resolution predictions. This method ensures that the model captures both fine details and global context effectively. The model operates on a high constant resolution throughout the process, enhancing its ability to generate detailed and coherent predictions across the entire image.
In this study, the DPT was directly applied to high-resolution RGB images collected from commercial apple orchards to generate detailed RGB-D (depth) data. The depth maps produced by DPT provided the 3D spatial context necessary for validating the pose estimation performed by the YOLO11 algorithm. Specifically, major axes of immature green fruits estimated by YOLO11 pose estimation model were compared against measurements obtained from 3D point clouds generated by DPT. The actual diameters of the fruitlets were also measured in the field using digital calipers to provide ground truth validation.
High-resolution RGB images were captured using specialized cameras mounted on a robotic platform as discussed before. These images were then preprocessed to enhance their quality, which were then used as inputs into the DPT model. The model processed these images through several stages, including patch extraction, transformer encoding, thresholding and feature map creation, resulting in RGB-D outputs. Subsequently, the model output was used to determine 3D coordinates of the key pints (opposite ends of the longitudinal axis of fruitlets) estimated by YOLO11 pose estimation model, which resulted in 3D poses and major axis lengths of the fruitlets. The estimated major axis length was then compared against the field measurements as discussed below.
@@ 3.4.2 Depth Anything V2
Depth Anything V2 is an advanced monocular depth estimation model designed to convert standard RGB images into depth-mapped (RGB-D) images Yang et al. (2024) . This model employs a fully convolutional network to interpret and transform visual data (on single RGB images) into spatial mappings. The architecture processes images through multiple layers to extract and refine features at various scales, capturing detailed and granular elements ideal for complex natural scenes. Utilizing an encoder-decoder network, the model compresses the image into a feature-rich format, and then expands these features to assign depth values across the image. Enhanced by skip connections that preserve spatial details, the model produces a dense depth map that accurately reflects the real-world geometry of the scene.
Similar to DPT, the Depth Anything V2 model was also utilized to estimate 3D locations of two key points of fruitlets (opposite ends of fruitlet’s major axis) estimated by the YOLO11 model as discussed before which is crucial before the thinning process in agricultural operations. As depicted in Figure 5b, the original RGB images (same images with 70 sample fruitlets mentioned before) collected using a consumer-grade camera were provided as inputs to Depth Anything V2 model. The images then went through processes including feature extraction, feature compression through an encoder to create a dense representation, and expanded representation with a decoder to generate detailed depth maps. The integration of skip connections helped preserve essential spatial details. The resulting depth maps provided comprehensive 3D spatial data, which was used to estimate the 3D coordinates of key points estimated by the YOLO11 model.
3.5 Field validation of estimated pose
RGB-D maps generated by two Vision Transformer models were used to extract resulting 3D point clouds (Figure 6)a. As discussed before, the image-based key points of fruitlets (two opposite ends of the fruitlet’s major axis) were converted into 3D coordinates in physical units using the point clouds generated by the transformer models as well as the point cloud captured by the RGB-D camera for comparison. Once 3D coordinates of the endpoints of the fruitlets were estimated, the distance between them was calculated (using a Euclidean distance formula) as the fruitlet size. To assess the accuracy of the estimated fruitlet sizes based on three different 3D point cloud datasets (two Vision Transformer generated and one sensor generated), ground truth fruit size (major axis length) was measured in the filed using a Vernier Caliper as illustrated in Figure 6b. Altogether, field measurements were collected for 70 randomly selected samples of immature green fruitlets. The accuracy assessment of fruitlet size estimation conducted with the in-field measurements served as a benchmark for validating the fruitlet poses estimated by the integrated YOLO11 and Vision Transformer models that were employed to estimate fruitlet key points and transform standard RGB images into detailed RGB-D data without additional depth information from the camera.